OpenStudy (anonymous):

Find the x- and y- intercepts for the graph of the following equation. 5x + 6y = 120

Parth (parthkohli):

You find the x-intercept by setting y = 0. Similarly for y-intercept, you set x = 0. Solve for the variables then.

OpenStudy (anonymous):

5x + 6y = 120 for finding x intercept we put y = 0 5x + 6*0 = 120 5x = 120 x = 120/5 = 24 so (24, 0) is the x intercept Similarly for y intercept we p[ut x = 0 5*0 + 6y = 120 6y = 120 y = 120/6 = 20 so y intercept is (0, 20)